OpenStudy (anonymous):

Identify the correct slope and y-intercept of the equation 5x - 3y = -6 slope = 5/3 ; y-intercept = (0, 2) slope = -5; y-intercept = (0, -6) slope =-3/5 ; y-intercept = (0, 2) slope =5/3 ; y-intercept = (0, -6)

OpenStudy (anonymous):

solve for y... can you do that first??

OpenStudy (anonymous):

and btw, y-intercept is when x=0... so can you eliminate some choices there?

OpenStudy (anonymous):

5 - 3y = -6 -5 -5 -3y = -11

OpenStudy (anonymous):

???

OpenStudy (anonymous):

what happened to the "x" ??

OpenStudy (anonymous):

slope intercept form equation is y=mx+c where m is you slope and c is your y intercept just put the equation in this form 5x - 3y = -6 5x+6=3x 5x/3+6/3=y (5/3)x+2=y y=(5/3)x+2 comparing with equation y=mx+c m=5/3 and c=2 or (0,2) so answer is 1